# From “Worse is Better” to Better: Lessons from a Mixed Methods Study of Ansible’s Challenges

### v1

@@ 7. Discussion
This section provides a set of actionable insights and discusses better IaC implications of our work.
@@ 7.1. Actionable Insights
This section proposes concrete actionable recommendations for improving the Ansible ecosystem. Each subsection addresses a key domain and references the findings from topic modeling and interviews to explain why these measures are timely and impactful.
7.1.1. Better Debugging Tools
@@
In our analysis of user’s Q&A data, reliability challenges — particularly “Error Handling” — were among the most frequently cited frustrations. Likewise, 10 interview participants described how diagnosing failures becomes particularly difficult when a playbook “gets stuck” with little to no logging feedback. Multiple interviewees lamented the absence of rollback-on-failure or built-in “step-by-step” debugging capabilities that might elucidate the exact point of failure. Although Ansible’s stateless design appeals to many users, it poses challenges for scenarios requiring continuous enforcement or “undo” functionality. While the Q&A data indicates that only a small subset of posts explicitly mention “dependency management” issues (see Figure 4), interviews reveal that participants struggle to remove outdated configurations in real-world environments.
Debugging support could include a “debug mode” generating structured, real-time logs, flagging potential syntax or connectivity issues earlier in the pipeline, and offering recommended fixes. To reconcile stateless simplicity with advanced rollback needs, we propose a configurable “soft state” mode that records minimal metadata—enough to revert undesired changes or compare current system states to a known baseline without imposing excessive overhead. Most debugging currently requires manual intervention, making troubleshooting time-consuming and error-prone. We propose the development of automated repair tools capable of identifying common misconfigurations, detecting inconsistencies in playbooks, and suggesting or even applying corrective actions automatically. Repair mechanisms exist in traditional software development, but their adoption in IaC environments remains limited (Weiss et al., 2017).
7.1.2. Higher-Level Language Concepts
@@
Although Ansible’s declarative approach and reliance on YAML/Jinja2 can simplify basic configuration tasks, our analysis of Q&A data found syntax issues with Jinja and other parts of Ansible were the main challenges of the Usability theme. Several interviewees (n = 10), mentioned the limitations in expressing complex logic. They described frustrations with YAML syntax quirks (especially for loops and conditionals), cumbersome variable handling, and a lack of “proper” language features for advanced use cases.
Addressing these challenges involves introducing higher-level language concepts. One possible approach is to bridge YAML with Python-based enhancements, allowing better support for loops, conditionals, or data processing without breaking Ansible’s declarative philosophy. This aligns with ongoing developments in Ansible’s SDK, which provides APIs for extending Ansible’s functionality using Python. Additionally, solutions like Pulumi (Pulumi, 2025) present an alternative paradigm where IaC is directly implemented in imperative programming languages rather than YAML-like configurations. While Ansible remains firmly declarative, allowing for optional hybrid workflows where advanced users can script automation logic in Python would reduce the need for error-prone YAML workarounds. We argue that the goal is not to replace YAML entirely but to augment it for scenarios that exceed basic configuration.
7.1.3. Improve Performance
@@
Our topic modeling analysis revealed a substantial volume of posts on slow execution (17.7% of the posts related to Performance and Scalability). In interviews, seven participants reported frustrations with performance bottlenecks, particularly when managing hundreds of hosts. They cited sequential task execution, agentless overhead, and limited parallelization features as key obstacles slowing down their workflows.
@@
Participants and our Q&A data analysis converge on the need for better parallelization and reduced overhead in large deployments. An alternative tool is Jetporch (Jetporch Contributors, 2025), a Rust-based Ansible alternative that aims to address performance limitations by avoiding Python (and its inherent execution overhead) and optimizing execution. Jetporch’s design ideas show key areas where Ansible could improve. Ansible should undergo refactoring to introduce parallelization improvements, optimize SSH connect, and enable partial execution in a more efficient manner.
7.1.4. Improve documentation
@@
Good documentation is one of the things that our participants liked about Ansible, so we argue that Ansible should capitalize on this fact by iterating over the current documentation, centralizing official resources, clarifying built-in vs. Galaxy modules, and providing version-specific references. Beyond traditional methods of documentation overhaul, generative AI tools could dynamically address knowledge gaps. For instance, LLMs could identify deprecated or outdated sections by continuously parsing official documentation and user-contributed content, and flagging inaccurate sections, automatically suggesting updates. Similarly, a tool like this could produce playbook snippets or configuration examples for less common scenarios on demand.
@@ 7.2. Future Work and Broad IaC Implications
Mixed-Methods Framework.
We observe differences between publicly discussed topics and the more subjective issues revealed through direct interviews. While thematic areas such as “Control Flow” or “Lack of Support” dominate forum posts, other challenges — like documentation gaps or configuration drift — appear more in interviews. For instance, information gaps and misleading advice may not prompt a direct question on a forum. Practitioners may not even realize they are facing an information gap, nor might they post about it explicitly. Instead, they continue troubleshooting, or they might consult colleagues offline. In interviews, however, participants openly discussed these intangible pain points — highlighting subjective experiences. We argue that qualitative insights should be used with quantitative or text-mining approaches to provide a holistic understanding of IaC adoption barriers. We provide a straightforward and replicable methodology that can be used for other major IaC platforms (e.g., Chef, Puppet) to yield more holistic, data-driven evaluations of challenges that practitioners face.
Balancing Usability and Scalability.
User experience seems to vary by expertise level. Beginners emphasize documentation clarity, onboarding simplicity, and accessible debugging tools. Meanwhile, advanced practitioners advocate for performance optimizations (e.g., concurrency) and advanced programming capabilities. This suggests that IaC tools should provide tiered user experiences — offering simplified workflows for beginners while exposing advanced performance optimizations and parallelization capabilities for experts. Striking this balance seems to be important to ensure long-term adoption across skill levels.
Documentation Gaps are a Universal Problem.
Finally, documentation gaps and information drift are persistent themes across IaC platforms, even in Ansible. While well-developed official documentation can lower the barrier to entry (Robillard, 2009; Aghajani et al., 2019), many tools struggle to keep pace with rapid feature expansion and a growing user base (Aghajani et al., 2020). Establishing centralized, consistently updated repositories of best practices and official modules is important for sustained usability.
@@ 8. Threats to Validity
Threats to Internal Validity.
A threat to internal validity arises from the use of LLMs through TopicGPT. LLMs are prone to hallucination and can generate unreliable results, potentially affecting the precision of our findings. To mitigate this threat, we perform manual validation in two stages: (1) during the generation of the topic list, and (2) during the assignment of topics to individual posts. Also, conducting semi-structured interviews introduces risks such as interviewer bias, where the phrasing or emphasis of questions might influence participant responses. To minimize this threat, we followed a consistent interview protocol and performed pilots before deploying the study.
Threats to External Validity.
A threat to external validity is that practitioners can discuss Ansible issues on alternative Q&A platforms that we did not analyze, potentially limiting the representativeness of our findings. To mitigate this, we collected posts from three distinct platforms. Another threat is that while some of our findings may generalize to other IaC tools, particularly configuration management tools, more research is needed to confirm their applicability. Future studies should explore other IaC tools to assess the generalizability of our results. Regarding the interviews, we recruited participants primarily through online channels, so our sample may not be representative of the broader Ansible user community. Self-selection bias may have resulted in participants who are more engaged with the tool or have stronger opinions about it. Furthermore, our sample’s distribution across expertise levels may not accurately reflect the actual distribution in the community.
@@ 9. Related Work
To identify challenges that practitioners face in IaC, Rahman et al. analyzed 2,758 Puppet-related questions from Stack Overflow (Rahman et al., 2018). Through qualitative analysis, the authors identified 16 major categories of questions. Many of these categories, including Syntax Error and Filesystem, are also observed in our study. While Rahman et al.’s study focuses on Puppet, our study focuses on Ansible, which is as of 2024 a more widely used technology for configuration management of services. Furthermore, Rahman et al. analyzed questions posted between January 2010 and December 2016, while our interviews and posts reflect the scenario as of 2024.
@@
Guerriero et al. conducted 44 semi-structured interviews with IaC practitioners to understand how practitioners adopt and develop IaC, the tools available within the IaC ecosystem along with their advantages and limitations, and the specific needs and challenges that practitioners face when working with IaC (Guerriero et al., 2019). Similarly to our study, they highlight issues such as difficulties with testing and debugging. However, we provide a more granular and in-depth analysis, as well as the identification of unique issues of Ansible, such as issues with loops and Jinja.
@@
Begoug et al. collected 52,692 questions and 64,078 responses from Stack Overflow to investigate challenges developers encounter with IaC (Begoug et al., 2023). The dataset includes questions related to Ansible, Pulumi, Terraform, CloudFormation, and other IaC technologies. The authors grouped questions into specific topics using the Latent Dirichlet Allocation (LDA) method optimized with a Genetic Algorithm (GA) for parameter fine-tuning. Our analysis of the questions of practitioners follows a similar approach. However, we collected data from two additional sources: Reddit and the Ansible forum. Additionally, instead of using LDA, we use TopicGPT for topic modeling, which not only provides interpretable topics, but also produces topics that align better with human categorizations than LDA (Pham et al., 2024). Finally, although we both found common topics such as File Management and Templating, due to our focus on Ansible, we found more granular and specific challenges faced by practitioners such as Jinja and AWX issues.
@@
To the best of our knowledge, the other only mixed methods study related to IaC challenges is the one conducted by Tanzil et al. (Tanzil et al., 2023). However, their study focuses on a broader scope, examining DevOps challenges in general, with IaC being one of the categories included. They used LDA to perform topic modeling on approximately 174,000 Stack Overflow posts, identifying 23 distinct topics. Then these topics were grouped into four categories, IaC being one of them, encompassing seven specific topics. Finally, the authors validated and extended their findings with a survey of 21 professional DevOps practitioners. Although there is some overlap in the findings between our study and that of Tanzil et al. (for example, both studies identify challenges related to File Management and Syntax), we focus exclusively on challenges related to Ansible, which provides a much narrower scope with different insights such as the issues on control flow, such as difficulties with loops and conditionals. We also employ a different topic modeling approach and focus on interviews over surveys.
@@ 10. Conclusion
This paper combined a large-scale quantitative analysis of 59,157 online Ansible discussions with 16 practitioner interviews to map out the most significant challenges, contextualize their real-world impact, and propose targeted solutions. We found issues related to control flow, performance bottlenecks, limited debugging capabilities, and inconsistent documentation, highlighting the tensions among usability, flexibility, and scalability within Ansible’s ecosystem. We offer concrete recommendations and show how Ansible can better evolve to serve novice users and large-scale deployments. These contributions, alongside broader reflections on community-driven knowledge, provide a replicable roadmap for assessing and improving IaC solutions.

### v2

@@ 7. Discussion
This section provides a set of actionable insights and discusses the implications of our findings for the IaC ecosystem.
@@ 7.1. Actionable Insights
This subsection proposes concrete, actionable recommendations for improving the Ansible ecosystem and the IaC ecosystem in general.
Debugging
@@
As noted in previous studies, debugging in IaC remains a significant challenge (Guerriero et al., 2019). In our analysis of users’ Q&A data, reliability challenges — particularly “Error Handling” — were among the most frequently cited frustrations. Likewise, 10 interview participants described how diagnosing failures becomes particularly difficult when a playbook “gets stuck” with little to no logging feedback. Multiple interviewees lamented the absence of rollback-on-failure or built-in “step-by-step” debugging capabilities that might elucidate the exact point of failure. Although Ansible’s stateless design appeals to many users, it poses challenges for scenarios requiring continuous enforcement or “undo” functionality.
For developers of IaC tools: In DevOps, where observability and rapid troubleshooting are essential, developers should make efforts to provide end users with comprehensive logging. Moreover, clear guidelines, or even mandatory standards, should be established for community-contributed modules to enforce consistent logging practices. IaC tools such as Ansible should develop automatic rollback mechanisms to ensure recovery from failures. Such mechanisms may rely on saving the system’s prior state or executing compensating operations to revert unsuccessful changes. Rollback support might enhance user confidence during IaC script testing and facilitate recovery from unexpected behavior.
For researchers: Participants emphasized the need for better debugging support in IaC tools as IaC tools lack the robust debugging utilities available in traditional languages like C and Python. Although Ansible offers a debugger (Red Hat, 2025a), practitioners reported difficulties and often resorted to shell scripts instead. Additionally, most debugging requires manual intervention, making troubleshooting time-consuming and prone to errors. We propose developing better debugging and automated repair tools that identify common misconfigurations, detect inconsistencies, and automatically suggest or apply fixes. While repair mechanisms exist in traditional software development, their adoption in IaC remains limited (Weiss et al., 2017).
Syntax
@@
Although Ansible’s declarative approach and reliance on YAML/Jinja2 can simplify basic configuration tasks, our analysis of Q&A data found syntax issues with YAML, Jinja and other parts of Ansible were the main challenges of the Usability theme. Additionally, loops and conditional executions accounted for a substantial portion of the reported difficulties. Several interviewees (n = 10) mentioned the limitations in expressing complex logic. They described frustrations with YAML syntax quirks (especially for loops and conditionals), cumbersome variable handling, and a lack of “proper” language features for advanced use cases.
For developers of IaC tools: Beyond the feedback obtained in our study, developers have consistently identified YAML as highly complex and error-prone. This issue is particularly concerning, given that YAML is used not only by Ansible but also by other IaC tools such as Kubernetes. To address this challenge, Asseldonk has proposed replacing YAML with simpler configuration languages, such as TOML (Tom Preston-Werner, 2025). However, to introduce higher-level language concepts, we must go beyond configuration languages. Solutions such as Pulumi (Pulumi, 2025) present an alternative paradigm where IaC is implemented in imperative programming languages rather than configuration languages such as YAML, while still maintaining declarative features. Another example is Nix (NixOS Contributors, 2025), which uses a functional programming language specifically designed for package definitions. Developers of IaC tools must carefully design or select the configuration language, as limitations in syntax and expressiveness can significantly influence the developers’ experience.
Performance
@@
Our topic modeling analysis revealed a substantial volume of posts on slow execution (17.7% of the posts related to Performance and Scalability). In interviews, seven participants reported frustrations with performance bottlenecks, particularly when managing hundreds of hosts. They cited sequential task execution, agentless overhead, and limited parallelization features as key obstacles slowing down their workflows. Participants and our Q&A data analysis converge on the need for better parallelization and reduced overhead in large deployments.
@@
For Ansible developers: Jetporch (Jetporch Contributors, 2025) is a Rust-based Ansible alternative that aims to address performance limitations by avoiding Python (and its inherent execution overhead) and optimizing execution. Jetporch’s design ideas show key areas where Ansible could improve. Ansible should undergo refactoring to introduce parallelization improvements, optimize SSH connections, and enable partial execution in a more efficient manner.
For developers of IaC tools: We recommend treating performance optimization as a priority. Developers should prioritize architectural improvements that enable efficient parallel execution across multiple hosts, optimize communication overhead (particularly for agentless tools relying on SSH), and provide configurable execution strategies for different deployment scales.
Documentation
@@
For Ansible developers: Good documentation is one of the things that our participants liked about Ansible, so we argue that Ansible should capitalize on this fact by improving its current documentation. Beyond traditional methods of documentation overhaul, generative AI tools could dynamically address knowledge gaps. LLMs could identify outdated sections by continuously parsing documentation and flagging inaccurate sections, automatically suggesting updates.
For developers of IaC tools: We recommend prioritizing comprehensive and well-maintained documentation. Developers should establish clear guidelines distinguishing official modules from community contributions, maintain version-specific documentation that is easily navigable, and implement regular audits to identify and update outdated content. Additionally, exploring AI-assisted documentation tools could help maintain accuracy at scale by automatically detecting deprecated information and generating contextual examples for edge cases.
@@ 7.2. Future Work and Broad IaC Implications
Mixed-Methods Framework. We observe differences between publicly discussed topics and the more subjective issues revealed through direct interviews. While thematic areas such as “Control Flow” or “Lack of Support” dominate forum posts, other challenges — like documentation gaps or configuration drift — appear more in interviews. For instance, information gaps and misleading advice may not prompt a direct question on a forum. Practitioners may not even realize they are facing an information gap, nor might they post about it explicitly. Instead, they continue troubleshooting, or they might consult colleagues offline. In interviews, however, participants openly discussed these intangible pain points — highlighting subjective experiences. We argue that qualitative insights should be used with quantitative or text-mining approaches to provide a holistic understanding of IaC adoption barriers. We provide a straightforward and replicable methodology that can be used for other major IaC platforms (e.g., Chef, Puppet) to yield more holistic, data-driven evaluations of challenges that practitioners face.
Balancing Usability and Scalability. User experience seems to vary by expertise level. Beginners emphasize documentation clarity, onboarding simplicity, and accessible debugging tools. Meanwhile, advanced practitioners advocate for performance optimizations (e.g., concurrency) and advanced programming capabilities. This suggests that IaC tools should provide tiered user experiences — offering simplified workflows for beginners while exposing advanced performance optimizations and parallelization capabilities for experts. Striking this balance seems to be important to ensure long-term adoption across skill levels.
Documentation Gaps are a Universal Problem. Finally, documentation gaps and information drift are well-documented challenges across software systems, and our findings confirm their presence in Ansible as well. While well-developed official documentation can lower the barrier to entry (Robillard, 2009; Aghajani et al., 2019), many tools struggle to keep pace with rapid feature expansion and a growing user base (Aghajani et al., 2020). Establishing centralized, consistently updated repositories of best practices and official modules is important for sustained usability.
@@ 8. Threats to Validity
Threats to Internal Validity. A threat arises from the use of LLMs through TopicGPT. LLMs are prone to hallucination and can generate unreliable results, potentially affecting the precision of our findings. To mitigate this, we perform manual validation in two stages: (1) during the generation of the topic list, and (2) during the assignment of topics to individual posts. Also, conducting semi-structured interviews introduces risks such as interviewer bias, where the phrasing or emphasis of questions might influence participant responses. To minimize this, we followed a consistent interview protocol and performed pilots before deploying the study.
Threats to External Validity. Practitioners can discuss Ansible issues on alternative Q&A platforms that we did not analyze, potentially limiting the representativeness of our findings. To mitigate this, we collected posts from three distinct platforms. Another threat is that while some of our findings may generalize to other IaC tools, particularly configuration management tools, more research is needed to confirm their applicability. Future studies should explore other IaC tools to assess the generalizability of our results. Regarding the interviews, we recruited participants primarily through online channels, so our sample may not be representative of the broader Ansible user community. Self-selection bias may have resulted in participants who are more engaged with the tool or have stronger opinions about it. Furthermore, our sample’s distribution across expertise levels may not accurately reflect the actual distribution in the community.
@@ 9. Related Work
To identify challenges that practitioners face in IaC, Rahman et al. analyzed 2,758 Puppet-related questions from Stack Overflow (Rahman et al., 2018). They used qualitative analysis to identify 16 major categories of questions. Many of these categories, including Syntax Error and Filesystem, are also observed in our study. While Rahman et al.’s study focuses on Puppet, our study focuses on Ansible, which, as of 2024, is a more widely used technology for configuration management of services. Furthermore, Rahman et al. analyzed questions posted between January 2010 and December 2016, while our interviews and posts reflect the scenario as of 2024.
@@
Guerriero et al. conducted 44 semi-structured interviews with IaC practitioners to examine adoption and development practices, available tools, and the needs and challenges of working with IaC (Guerriero et al., 2019). Similarly to our study, they highlight issues such as difficulties with testing and debugging. However, we provide a more granular and in-depth analysis, as well as the identification of unique issues of Ansible, such as issues with loops and Jinja.
@@
Begoug et al. collected 52,692 questions and 64,078 responses from Stack Overflow to investigate challenges developers encounter with IaC (Begoug et al., 2023). The dataset includes questions related to Ansible, Pulumi, Terraform, CloudFormation, and other IaC technologies. The authors grouped questions into topics using the Latent Dirichlet Allocation (LDA) method optimized with a Genetic Algorithm (GA) for parameter fine-tuning. Our analysis of practitioners’ questions follows a similar approach. However, we collected data from two additional sources: Reddit and the Ansible forum. Additionally, instead of using LDA, we use TopicGPT, which not only provides interpretable topics but also produces topics that align better with human categorizations (Pham et al., 2024). We both found common topics such as File Management and Templating, but we found additional challenges faced by practitioners such as Jinja and AWX issues. Another key distinction between our study and those described above is that while they focus either on interviews or on questions posted by practitioners, our study employs a mixed-method approach that integrates data from both sources.
@@
To the best of our knowledge, the only other mixed methods study related to IaC challenges is the one conducted by Tanzil et al. (Tanzil et al., 2023). However, their study is broader, addressing DevOps challenges in general, with IaC as just one category. They used LDA to perform topic modeling on approximately 174,000 Stack Overflow posts, identifying 23 distinct topics. These were then grouped into four categories, with IaC being one of them, encompassing seven specific subtopics. Finally, the authors validated and extended their findings with a survey of 21 professional DevOps practitioners. Although there is some overlap in the findings between our study and that of Tanzil et al. (e.g., both identify challenges related to File Management and Syntax), we focus exclusively on challenges related to Ansible and IaC, which provides a more focused scope with different insights such as difficulties with loops and conditionals. We also employ a different topic modeling approach and focus on interviews over surveys.
@@ 10. Conclusion
We combined a large-scale quantitative analysis of 59,157 online Ansible discussions with 20 practitioner interviews to identify key challenges, contextualize their real-world impact, and propose solutions. We found issues related to control flow, performance bottlenecks, limited debugging capabilities, and inconsistent documentation, highlighting the tensions among usability, flexibility, and scalability within Ansible’s ecosystem. We offer concrete recommendations and show how Ansible can better evolve to serve novice users and large-scale deployments. These contributions, alongside broader reflections on community-driven knowledge, provide a replicable roadmap for assessing and improving IaC solutions.